OpenStudy (anonymous):

2x-3y=-1 Y=x-1 solve by substitution help ASAP

OpenStudy (texaschic101):

y = x - 1 now sub x - 1 in for y in the other equation 2x - 3y = -1 2x - 3(x - 1) = -1 2x - 3x + 3 = -1 2x - 3x = -1 - 3 -x = -4 x = 4 now sub 4 in for x in either equation to find y y = x - 1 y = 4 - 1 y = 3 check... 2x - 3y = -1 2(4) - 3(3) = -1 8 - 9 = -1 -1 = -1 (correct) x = 4 and y = 3